3x-2.5y=7.125
2.5x-3y=7.3125
I have to solve by using the addition method. Am I supposed to convert these? If so, I have no idea how.

Answer by checkley77(12844)

3x-2.5y=7.125 multiply by -2.5
2.5x-3y=7.3125 multiply by 3 & add
-7.5x+6.25=-17.8125
7.5x-9y=21.9375
--------------------------
-3y=-4.125
y=-4.125/-3
y=1.375 ans.
3x-2.5*1.375=7.3125
3x-3.4375=7.3125
3x=7.3125+3.4375
3x=10.75
x=10.75/3
x=3.5833 ans.
Proof:
2.5*3.5833-3*1.375=7.3125
8.958-4.125=4.833
4.833=4.833